OpenStudy (mysticfury):

Enter an equation that can be used to solve this problem, where x represents the number of hours that the first plane had been traveling. Two planes left the same airport traveling in opposite directions. The first plane left the airport at 2 P.M. traveling at an average rate of 500 mph. The second plane left the airport at 4 P.M. traveling at an average rate of 550 mph. When will the planes be 3100 miles apart?

OpenStudy (mysticfury):

@Awolflover1 @AloneS @pooja195 @agent0smith

OpenStudy (mysticfury):

500x - 550x = 3100 this is what i have currently but, its not right and idk where to go from here

OpenStudy (mysticfury):

@hartnn @zepdrix @Nnesha

OpenStudy (mysticfury):

@mathstudent55

OpenStudy (kittiwitti1):

well I think they would not be both \(x\) as they left at different times...

Awolflover1 (awolflover1):

500 x - 1100 + 550 x = 3100 ? I'm confused

OpenStudy (mysticfury):

i was thinking more of 550(x - 2) - 500x = 3100?

OpenStudy (kittiwitti1):

That might work, yes.

Awolflover1 (awolflover1):

yeah

OpenStudy (mysticfury):

hmm

OpenStudy (mysticfury):

@kidrah69 @whpalmer4 @YanaSidlinskiy @welshfella

OpenStudy (kittiwitti1):

Hm? Are you still confused?

OpenStudy (mysticfury):

yeah

OpenStudy (kittiwitti1):

Where at? o:

OpenStudy (mysticfury):

im confused on how to make the equation

OpenStudy (kittiwitti1):

I think you were going in the right direction.

OpenStudy (welshfella):

for the first 2 hours the first plane travels 1000 miles ( as it travles at 500 mph) At 4 pm they are travelling apart at a rate of 500 + 550 = 1050 mph

OpenStudy (mysticfury):

ok..

OpenStudy (welshfella):

so they travel another 2100 miles after the second plane takes off. During this journey the first plane will travel for (x - 2) hours so our equation is 1050( x - 2) = 2100 solve for x

OpenStudy (mysticfury):

okay then . it dosent want you to solve it just to put an equation

OpenStudy (welshfella):

yes

OpenStudy (mysticfury):

wait 2100 ? why 2100?

OpenStudy (welshfella):

3100 - 1000

OpenStudy (mysticfury):

oh so 1050(x - 2) = 2100 is the equation?

OpenStudy (welshfella):

the planes are 1000 miles apart when the second plane takes off.

OpenStudy (mysticfury):

but since the second plane is at a faster speed wouldnt that mean the planes get closer together instead of farther away?

OpenStudy (welshfella):

yes where x = number of hours the first plane travels he has already been 2 hours in the sir before the second takes off. Thats why its (x - 2) hours for the journey of 2100 miles

OpenStudy (welshfella):

No - they are travelling in opposite directions.

OpenStudy (mysticfury):

ohhhhhh

OpenStudy (mysticfury):

okay i get it now , since they are traveling in opposite directions, i forgot about that

OpenStudy (welshfella):

so they travel apart at the rate of the sum of their speeds 500 + 550

OpenStudy (mysticfury):

yep yep, okay i thought they were going in the same direction XD thanks a lot!

OpenStudy (welshfella):

yw

OpenStudy (agent0smith):

A diagram would've helped you make sense of the opposite directions part.

OpenStudy (agent0smith):

"okay i thought they were going in the same direction XD" or just reading the question clearly lol